If you put up $26784 today in exchange for a 9.3 percent, 17-year annuity, what will the annual cash flow be? (Round time value factors to 6 decimal places and final answer to the nearest dollar amount. Omit the "$" sign and commas in your response. For example, $123,456 should be entered as 123456.)
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

26784=Annuity[1-(1.093)^-17]/0.093

26784=Annuity*8.381455

Annuity=26784/8.381455

which is equal to

=3196(Approx).
Answer #2

To calculate the annual cash flow for a 9.3 percent, 17-year annuity with a present value of $26,784, we can use the formula for the annuity payment:

Annual cash flow = Present value × (Annuity factor)

Annuity factor can be calculated using the formula:

Annuity factor = (1 - (1 + r)^(-n)) / r

where r is the interest rate per period (9.3% = 0.093 as a decimal) and n is the number of periods (17 years).

Let's plug in the values and calculate the annual cash flow:

r = 0.093 (interest rate per period) n = 17 (number of periods) Present value = $26,784

Annuity factor = (1 - (1 + 0.093)^(-17)) / 0.093 Annuity factor ≈ 11.139106

Annual cash flow = $26,784 × 11.139106 ≈ $298,087.91

The annual cash flow will be approximately $298,087.91.

answered by: Hydra Master